Suppose that 60 percent of women who purchase over the counter pregnancy tests are actually pregnant. Call this event B.

1. what is Bc?

2. the probability that the pregnancy test id positive given that the person is pregnant is 0.96 (true positive). what is the probability that the pregnancy test is negative given that the person is pregnant? (False negative).

3. the probability that the pregnancy test is positive given that the person is not pregnant is 0.01 (false positive). what is the probability that the pregnancy test is negative given that the person is not pregnant? (true negative).

4. what is the probability that the person is pregnant given the pregnancy test is positive? P(B/A).
Answer #1

1)P(Bc) =1-P(B) =1-0.60 =0.40

2)

probability that the pregnancy test is negative given that the person is pregnant =1-0.96 =0.04

3)

probability that the pregnancy test is negative given that the person is not pregnant =1-0.01 =0.99

4)

P(B|A) =0.6*0.96/(0.6*0.96+0.4*0.01)=0.9931